OpenStudy (anonymous):

it gives me a right triangle prism, i guess, the height is 6.4 the base is 2 and the line at the top of the hypotenuse is 2.9. find the surface area

OpenStudy (anonymous):

I'm going to send a pic. Tell me if this is what you have.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

6.4, not 6...sorry

OpenStudy (anonymous):

how do you draw something

OpenStudy (anonymous):

I just go into 'paint' in windows, then save it on the desktop, then click 'attach file' below, attach and post.

OpenStudy (anonymous):

If it's what I think it is, it isn't 'hard', just finicky.

OpenStudy (anonymous):

It comes down to the same principle we used to solve the last problem. You have faces that are shapes you already know the area of. You just break the surface area down into those shapes. Are you drawing something?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

nothing shows up

OpenStudy (anonymous):

Ah, okay, yes, something did come through - I just had to magnify your atom of a picture, lol!

OpenStudy (anonymous):

how do u magnify and it was pretty big when i made it

OpenStudy (anonymous):

I'm thinking the hardest part for you is the triangular face?

OpenStudy (anonymous):

I use Google Chrome, and it opens in a new page with a little magnifying icon.

OpenStudy (anonymous):

Let's do the triangle faces first.

OpenStudy (anonymous):

where is it at when u clicked it in the center

OpenStudy (anonymous):

and ok

OpenStudy (anonymous):

You have two right-angled triangles. These have a height of 6.4 and base 2. The area of any triangle is \[\frac{1}{2}(base)(perpendicular.height)\]Luckily, perpendicular height is the height of a right-angled triangle. So your area is\[\frac{1}{2} \times 2 \times 6.4=6.4\]You have two of these, one on the top, one on the bottom.

OpenStudy (anonymous):

So the area of the triangular faces is 2 x 6.4 = 12.8. Put this aside and tackle the faces on the side.

OpenStudy (anonymous):

These are just rectangles. Each has dimensions 6.4 x 2.9 = 18.56. Again, you have two of these, so your area for the side-on rectangular faces is 2 x 18.56 = 37.12.

OpenStudy (anonymous):

Now all you gave left is the base. It is rectangular and has dimensions, 2 x 2.9 = 5.8. Can you see how the side length of the base is the same as the length of the prism at the top (where you've written 2.9)?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

then just add them all up?

OpenStudy (anonymous):

The total area is just the sum of these separate areas:\[12.8 + 18.56+5.8=37.16\]square units.

OpenStudy (anonymous):

You got it.

OpenStudy (anonymous):

youre freakin smart man

OpenStudy (anonymous):

You picked it up, so you're doing alright too ;)